OpenStudy (anonymous):

The temperature at a mountain base camp was −4 degrees Celsius on Saturday. Sunday morning, the temperature was 2 degrees Celsius lower than it was on Saturday. By Sunday evening, the temperature was 1 degree Celsius lower than it was that morning. What was the temperature at the base camp Sunday evening?

OpenStudy (anonymous):

need help right away

OpenStudy (anonymous):

wiil medal and fan

OpenStudy (dancergirl45):

hold on

OpenStudy (anonymous):

Your not supposed to give answers out on here, just a heads up, you should explain how to get the answer but not actually give it out.

OpenStudy (michele_laino):

Sunday morning the temperature is: \(-4-2=...?\)

OpenStudy (dancergirl45):

Well it would be -7 because -5-2-1=-7

OpenStudy (michele_laino):

better is: \(-4-2-1=...?\)

OpenStudy (dancergirl45):

sorry -4

OpenStudy (anonymous):

how did you get that? @dancergirl45